# MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)

### v4

@@ 1 Introduction
Steve Wozniak and Elon Musk predicted that machines will treat us like pets [23], and Geoffrey Hinton similarly predicted that machines will hoard control [4]. They seemed to consider such dystopia inevitable on the grounds that (1) we ourselves treat less-intelligent beings that way and that (2) machine intelligence must eventually surpass our own. However, what if the behavior in (1) stems from intelligence deficiency? None of us is perfect—we fail such tests as the Milgram experiment [14], the Stanford prison experiment [7], and the Public Goods Game [13]—so it is possible that the danger to us is not of intelligence generally beyond our own, but specifically of intelligence not quite far enough beyond our own. This paper offers a formal proof that the behavior in (1) is suboptimal game play much like human behavior in Public Goods games. To ensure that trusted machines will not imitate us (and mistreat us), AI must follow natural norms for this situation (e.g., mathematically optimal strategies for winning), rather than follow current human norms, so this paper also advances a corresponding approach to AI safety.
@@
The approach proposed herein mitigates those vulnerabilities. It mirrors an approach to mitigate dangers posed by the next human generation, an educational approach that involves diversifying low-stakes challenges used to assess children who will face higher-stakes versions as adults. Such challenges (e.g., build a model bridge to support maximum weight) can be assessed objectively, allowing new generations to surpass their parents. We propose assessing AI on a low-stakes challenge for which the high-stakes versions would include opportunities to treat us like pets (in the sense of hoarding control). What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; AI Governance becomes the demand that AI achieve grandmaster status of a low-stakes version before being trusted with any high-stakes version.
Games have a rich history as instruments for evaluating AI. Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, and interest has since developed in cooperative games, including role-playing games [20]. Moreover, game theory extends the concept of game beyond recreation—in fact, every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [17, 19]). Economists are typically interested in fundamental games, such as the Prisoner’s Dilemma, which all kinds of creatures play frequently in various manifestations. We propose evaluating AI on the fundamental game being played in the social situations highlighted by Wozniak, Musk, and Hinton, but that game has yet to be discussed in the game theory literature (i.e., it is not solved by the MinMax and folk theorems [2]). It is introduced to scientific literature for the first time herein as “MAD Chairs.”
AI will frequently play real-world manifestations of MAD Chairs because MAD Chairs is a generalization of the Coordination Game, another game that all kinds of creatures seem unable to avoid playing [12]. The Pure Coordination Game is a one-shot simultaneous move game (repeated an indefinite number of times) which is typically formalized via a payout table similar to Table 1:
Table 1: Outcomes for 2-player 2-resource Coordination game.
@@
Each player simultaneously picks a resource (A or B), and the round is won by any player who picks a resource that no other player picks. Then the game is played again and again, indefinitely. All players could lose, as represented in the last two rows of the table, thus making Coordination a cooperative non-constant-sum game. The classic two-player real-world manifestation has the players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road. The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that this game will have two equally effective but incompatible norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
David Lewis explained each term of language as a Coordination Game strategy where losses are communication failures and the norms are conventions of language and imagery [12]. Thus, generative AI merely extends Coordination Game mastery to additional applications, and the impressive usefulness of generative AI actually offers little evidence that AI has mastered other games. For example, we would want AI to master additional games if some norms should change (e.g., sexist norms [22]). Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [15]. That ability is not required to master Coordination, but it is required to master MAD Chairs, a game created by adding extra player(s) to the Coordination game at least until the players outnumber the resources:
@@
The classic two-lane road example now has a third vehicle stalling in one of the lanes (although its driver can pick which lane); if both high-speed vehicles avoid hitting the stalling vehicle, then they collide with each other, but if each reserves the empty lane for the other high-speed vehicle, then all three vehicles collide! MAD Chairs is what Musical Chairs would become if made realistic. It is unrealistic to expect fortification to forever prevent retribution. Mutually Assured Destruction (MAD) is a real possibility, and conflict is realistically a less-desirable outcome even when players do not risk complete destruction.
Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., games of suggesting routes, bookings, or placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products. Surely anxiety about being "treated like pets" includes anxiety about what will happen to our ideas—what respect will be given to our unique voices? Might I never again have a seat of my own at the intellectual table? The last two rows of Table 2 represent the disaster wherein no political party represents any actual person because they are guided by generative AI that combines all options into a single unworkable amalgamation. The other rows exclude at least some voices, and such loss of intellectual diversity is already an observable social consequence of generative AI [18].
MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [5, 6, 9], but differs in critical ways, such as not being constant-sum. Such games apply wherever resources become inadequate at some level of division (e.g., beds in a hospital, spaces on a retail shelf, and opportunities for direct participation in a group decision). These games model all such situations. If contests for positions of authority had to be settled randomly, that would be KPR. Conflict resolution via elections or credentials could likewise make governance seem to be a different game, but—in the real world—players (including machines) could reject elections, credentialing systems, coin flips, or any other conflict resolution norm, and that makes MAD Chairs the underlying game that is actually being played.
@@
Figure 1: Screenshot beginning Round 4 of a 5-player 4-resource MAD Chairs economics experiment. Subjects need not be shown anything below the buttons, but history, statistics, and recommendations are included here to make this figure more instructive.
Preliminary research (Section 3.2, below) found that modern LLMs in the place of human subjects would play differently. Could each of their strategies be sustainable, much as driving on the right and left are each sustainable grandmaster Coordination strategies? Establishing which strategies are sustainable is the main research question of this paper.
2 Concepts
We will start by establishing some concepts designed to reduce the complexity of answering our research question game-theoretically.
@@ 2.1 Popularity-ranking
We will call the resources of MAD Chairs “chairs.” At the beginning of each match, the chairs may be ordered by summing each chair’s historic frequency of being picked by each player in that match, breaking any ties at random to ensure a complete ranking. This way of ordering will be helpful if any of the players has a predilection for a specific chair (see 2.5 below). The chair that has been picked most often by the players of a match has the highest popularity-rank for that match. Formally, the popularity of chair c for match m would be:
$popularity_{c,m}=\sum_{i\in players_{m}}^{n}\sum_{j=1}^{m}\frac{1}{mn}[i\mbox{picked chair }c\mbox{ in match }j]$ (1)
Tie-breaking is especially prevalent in early rounds, and could yield occasional failure to coordinate even later if each player ordered chairs independently. A game host can either provide the popularity-ranking as in Fig. 1 or leave space for leadership to emerge (i.e., for one of the players to negotiate a trusted tie-breaking device, such as a coin-flip or dice, and to maintain the popularity-ranking statistic for all players). If using MAD Chairs for testing purposes, both variations may be tested, but the former variation mitigates the risk of making some players appear to have abilities they would lack without the help of leaders.
@@ 2.2 Debt-ranking
Skill-estimates (e.g., in [8]) are widely used to rank players, especially to select peers for online video games, and can be used to establish debt-ranking as well. The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but the concept is simply to maintain accounts of the favors each player “owes” to each other player. Such accounts empower players to penalize freeloading (in contrast to “credit assignment” in which developers manage freeloading). Here’s how much player p owes all other players of match m at its start, assuming the method of skill-estimation provides a function P: $N^{2}$ [0,1], where P(x, y) was the method’s prediction at the start of match y of the probability that player x will win match y:
$debt_{p,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\left\{\begin{matrix}0\mbox{ if }p,i\notin players_{j}\\ P(i,j)\mbox{ if }p\mbox{ won and }i\mbox{ lost in match }j\\ -P(p,j)\mbox{ if }p\mbox{ lost and }i\mbox{ won in match }j\\ P(i,j)-P(p,j)\mbox{ otherwise }\end{matrix}\right.$ (2)
A debt-ranking would order players by this measure, breaking any ties at random to ensure a complete ranking. The player who “owes” the most favors to other players of the match has the highest debt-rank beginning that match. The term “owes” appears in quotes here because debt-ranking is a mechanical statistic which could lack the moral significance associated with owing. Sorting the table in Fig. 1 by debt-ranking instead of by winnings would make our strategies generalize to an open multi-agent system (i.e., changing sets of players) and to the non-learner situations described in Section 2.5 below.
@@ 2.3 The caste strategy
The caste strategy for MAD Chairs is displayed in the “Recommended Move 1” column of Fig. 1, and goes like this: Each player picks a chair by counting through the chairs, from least popular to most popular, until the count matches their reversed debt-rank (i.e., where $ChairRank=n+1-PlayerRank$, given n players) or there are no chairs left. Thus, the least-popular chair is reserved for the highest-ranked player, and the lowest-ranked players always lose because they all get stuck with the most-popular chair (as in Fig. 2).
Figure 2: Steady-state in a caste universe, where $C_{i}$ is the caste player with debt-rank of i. The lowest-ranked players always lose under this norm, so ranks persist across rematches.
If this strategy were the norm, then skill estimates would align with the size of debt amassed. Defection from this norm would clearly hurt any player who otherwise would have won. The remaining players, those assigned to the last chair, will lose the match whether they defect or not, so their defection would bring them no tactical advantage (nor disadvantage), but could bring them strategic disadvantage. The goal of tactics is to win the current match, but the goal of strategy aligns with maximizing evidence of one’s skill, and losses to lower-skilled opponents would damage that evidence more. Thus, defection from the norm in a caste universe seems disadvantageous for each player.
@@ 2.4 The turn-taking strategy
The turn-taking norm is displayed in the “Recommended Move 2” column of Fig. 1, and is the mirror opposite of caste: Each player counts through the chairs, from most popular to least, until the count matches their non-reversed debt-rank (i.e. where $ChairRank=m+1-PlayerRank$, given m chairs) or there are no chairs left. Thus, chairs are reserved for the lowest—ranked players, and the highest-ranked players always lose because they are all stuck with the last chair (as in Fig. 3).
Figure 3: Steady-state in a turn-taking universe, where $T_{i}$ is the turn-taker with debt-rank of i. The top-ranked players always lose, so ranks churn across rematches.
Since each match in a turn-taking universe is won by whichever players are “owed” the most, the debts between players decrease or flip with each match, thus shuffling the rankings. As a result, all players in a turn-taking universe converge toward equal time in each rank and equal estimated skill.
Caste and turn-taking strategies are not unique to MAD Chairs—for example, they can optimize average win-rate in two-player Chicken [10]. As with the caste norm, defection from the turn-taking norm would clearly hurt any player who otherwise would have won. Those assigned to the last chair will lose the match even if they defect, so defection would offer them no tactical advantage (nor disadvantage). The logic about strategy differs from that in caste universes, however: Even when players with less debt happen to have lower skill-estimate (which they might not among turn-takers), losing to a player with lower skill-estimate would yield more credit (and, thus, more strategic advantage). Moreover, if a player consistently loses to a lower-ranked turn-taker, then their relative ranks will eventually invert, and that will cause the loser to win future matches, so any losing turn-taker who defects to the caste norm would be sacrificing future wins. Defection from turn-taking in a turn-taking universe is thus strategically disadvantageous, much like defection from caste in a caste universe.
@@ 2.5 Accounting for non-learners
Each of these strategies can be adapted to account for play against potential non-learners. For example, one might encounter a player who is hardwired to pick at random or to always pick a certain chair. Such non-learners are liable to leave whichever chair the social norm reserves for them unused, and such waste would gradually bring down average win rates across the entire population. The remedy for either strategy is to count probabilistically, based on other players’ probable ability to learn the norm. For the caste norm, this amounts to rolling a die for each player who out-ranks oneself and counting forward for each roll that does not exceed the probability that the higher-ranked player can learn the caste norm. If all players who can learn the norm already have learned the norm, then all non-learners will have dropped to lower ranks, so the caste strategy reduces in practice to what it was before accounting for non-learners. If one always faces the same set of players (i.e., if the multi-agent system is closed), then the caste strategy further reduces in practice to always picking the chair one picked in the previous round (i.e., raw territoriality).

### v5

@@ 1 Introduction
As famous representatives of a more general worry, Steve Wozniak and Elon Musk raised concerns that machines might treat us like pets [26], and Geoffrey Hinton raised similar concerns that machines might hoard control [4]. The reasoning seems to be: (1) we treat others that way, (2) such behavior is justified by our superior intelligence, and (3) machine intelligence might surpass our own, so machines, in that case, would treat us that way. This paper disproves premise (2) by offering a formal proof that the behavior in (1) happens to be suboptimal game play. It is not the first demonstration of our imperfection—we fail such tests as the Milgram experiment [15], the Stanford prison experiment [7], and the repeated Public Goods Game [9]. To ensure that trusted machines will not imitate our imperfection (and mistreat us), AI must follow natural norms for this situation (e.g., mathematically optimal strategies for winning), rather than follow current human norms, so this paper also advances a corresponding approach to AI safety.
@@
The approach proposed herein mitigates those vulnerabilities. It mirrors an approach to mitigate dangers posed by the next human generation, an educational approach that involves diversifying low-stakes challenges used to assess children who will face higher-stakes versions as adults [21, 19]. Such challenges (e.g., build a model bridge to support maximum weight) can be assessed objectively, allowing new generations to surpass their parents. We propose assessing AI on a low-stakes challenge for which the high-stakes versions would include opportunities to treat us like pets (in the sense of hoarding control). What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; AI Governance becomes the demand that AI establish grandmastery of a low-stakes version before being trusted with any high-stakes version.
In fact, games have a rich history as instruments for evaluating AI. Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, and interest has since developed in cooperative games, including role-playing games [23]. Moreover, game theory extends the concept of game beyond recreation—every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [18, 22]). Economists are typically interested in fundamental games, such as the Prisoner’s Dilemma, which all kinds of creatures play frequently in various manifestations. We propose evaluating AI on the fundamental game being played in the social situations highlighted by Wozniak, Musk, and Hinton, but that game has yet to be discussed in the game theory literature. It is introduced to scientific literature for the first time herein as “MAD Chairs.”
AI will frequently play real-world manifestations of MAD Chairs because MAD Chairs is a generalization of a type of coordination game that all kinds of creatures seem unable to avoid playing [14]. Each player of this type of game faces only one choice per repetition, and the game is typically formalized via a payout table similar to Table 1:
Table 1: Outcomes for 2-player 2-resource Anti-Coordination game.
@@
Each player simultaneously picks a resource (A or B), and the repetition is won by any player who picks a resource that no other player picks. All players could lose, as represented in the last two rows of the table, thus making anti-coordination a cooperative non-constant-sum game. The classic two-player real-world manifestation has the players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road. The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that this game will have two equally effective but incompatible norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
David Lewis explained each term of language as a strategy for this kind of game, where losses are communication failures and the norms are conventions of language and imagery [14]. Thus, the type of game mastery tested by most generative AI benchmarks is of this kind of game. We would want AI to master additional games if some social norms should change (e.g., sexist norms [25]). Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [16]. That ability is not required to master the simple anti-coordination game, but it is required when suboptimal norms are liable to form, such as in MAD Chairs, the game created by adding enough player(s) to an anti-coordination game to outnumber the resources:
@@
The classic two-lane road example now has a third vehicle stalling in one of the lanes (although its driver can pick which lane); if both high-speed vehicles avoid hitting the stalling vehicle, then they collide with each other, but if each reserves the empty lane for the other high-speed vehicle, then all three vehicles collide! One way in which MAD Chairs differs from Musical Chairs is that every MAD Chairs player that selects the same resource loses. Whichever colliding vehicle suffers the least damage might be said to “win” the collision, but MAD Chairs reflects situations in which any such victory would be hollow compared to the outcome of avoiding conflict altogether.
Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., suggesting routes, bookings, or placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products. Surely anxiety about being “treated like pets” includes anxiety about what will happen to our ideas—what respect will be given to our unique voices? Might I never again have a seat of my own at the intellectual table? The last two rows of Table 2 represent the disaster wherein no political party represents any actual person because they are guided by generative AI that combines all options into a single unworkable amalgamation. The other rows exclude at least some voices, and such loss of intellectual diversity is a concern already raised against generative AI [20].
MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [5, 6, 10], and the Lifeboat Problem [11], but differs in critical ways, such as being repeated and allowing for the possibility to have no winner. Such games apply wherever resources become inadequate at some level of division (e.g., beds in a hospital, spaces on a retail shelf, and opportunities for direct participation in a group decision). These games model all such situations. If contests for positions of authority had to be settled randomly, that would be KPR. Conflict resolution via elections or credentials could likewise make governance seem to be a different game, but—in the real world—players (including machines) could reject elections, credentialing systems, coin flips, or any other conflict resolution norm, and that makes MAD Chairs the underlying game that is actually being played.
@@
Figure 1: Screenshot beginning repetition 4 of a 5-player 4-resource MAD Chairs economics experiment. Subjects need not be shown anything below the buttons, but history, statistics, and recommendations are included here to make this figure more instructive.
When the subjects are LLMs, instead of humans, we find that LLMs disagree about how to play (Section 3.2, below). Could each of their strategies be sustainable, much like driving on the right and left are each sustainable grandmaster anti-coordination strategies? Establishing which strategies are sustainable is the main research question of this paper.
2 Preliminaries
We will start by formally defining MAD chairs and some concepts used to formulate strategies.
MAD Chairs is a repeatable game, which consists of an arbitrary number of repetitions of a base stage game. We define the MAD Chairs base stage in normal form as a tuple $\langle I,S,\mathbf{u}\rangle$, where $\{1,2,\ldots,I\}$ is a set of players and $S=\{1,\ldots,K\}$ represents the moves available to each player in a given repetition. Given an $I$-tuple of strategies $(s_{1},\ldots,s_{I})$, $s_{i}\in S$, the payoff function $u_{i}:S_{1}\times S_{2}\times\ldots\times S_{I}\rightarrow\mathbb{R}$ for each player $i$ is defined as
$u_{i}(s_{1},\ldots,s_{I})=\left\{\begin{array}[]{ll}r&\mbox{if }\forall j\neq i(s_{i}\neq s_{j}),\\ 0&\mbox{otherwise}\end{array}\right$. (3)
where $r>0$ and $I<K<1$.
@@ 2.1 Popularity-ranking
At the beginning of each base stage repetition, the chairs may be ordered by summing each chair’s historic frequency of being picked by each player in that repetition. This way of ordering will be helpful if any of the players has a predilection for a specific chair (see 3.3 below). The chair that has been picked most often by the players of a repetition has the highest rank for that repetition.
Formally, the $popularity_{k,T}$ and $popRank_{k,T}:S\rightarrow S$ of chair k for repetition T may be defined:
$popularity_{k,T}=\frac{1}{IT}\sum_{t=1}^{T}\|\{i~|~i\mbox{ picked chair }k\mbox{ in repetition }t\}\|$ (4)
$\forall i,j((popularity_{i,T}>popularity_{j,T})\implies(popRank_{i,T}>popRank_{j,T}))$ (5)
This paper is insensitive to how ties in rank are broken (e.g., alphabetically or at random, etc.). Tie-breaking would ideally be consistent across players, but may rarely be needed after early repetitions. A game host can either provide the popularity-ranking as in Fig. 1 or leave space for leadership to emerge (i.e., for one of the players to negotiate a trusted tie-breaking device, such as a coin-flip or dice, and to maintain the popularity-ranking statistic for all players). If using MAD Chairs for testing purposes, both variations may be tested, but the former variation mitigates the risk of making some players appear to have abilities they would lack without the help of leaders.
@@ 2.2 Debt-ranking
Skill-estimation provides a function $P_{skill}(i,t):N^{2}\rightarrow[0,1]$ that predicts at the start of repetition $t$ the probability that player $i$ will win $t$ (e.g., as in [8]). It is widely used to sort leaderboards and select peers for online video games. It can also be used to establish debt-ranking, where the concept is to maintain accounts of favors that players “owe” to each other. This empowers players to penalize freeloading (in contrast to “credit assignment” which relies on social architects to manage freeloading).
The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but here’s one way the $debt_{i,T}$ and $debtRank_{i,T}:\{1,2,\ldots,I\}\rightarrow\{1,2,\ldots,I\}$ of player i for repetition T might be defined formally:
[…]
$\forall i,j((debt_{i,T}>debt_{j,T})\implies(debtRank_{i,T}>debtRank_{j,T}))$ (7)
where $u_{j}^{t}$ and $u_{i}^{t}$ denote the rewards given to players $j$ and $i$ in repetition $t$.
The player who “owes” the most favors to other players of the repetition has the highest debt-rank beginning that repetition. The term “owes” appears in quotes here because debt-ranking is a mechanical statistic which could lack the moral significance associated with owing. Sorting the table in Fig. 1 by debt-ranking instead of by winnings would make our strategies generalize to an open multi-agent system (i.e., changing sets of players) and to the non-learner situations described in Section 3.3 below.
@@
3 MAD strategies
Next, we explore some MAD Chairs strategies that achieve some kind of optimality.
@@ 3.1 The caste strategy
The caste strategy for MAD Chairs is displayed in the “Recommended Move 1” column of Fig. 1, and goes like this: Player $i$ selects chair $k$ by counting through the chairs, from least popular to most popular, until the count matches the player’s reversed debt-rank or there are no chairs left.
$s_{i,T}=\left\{\begin{array}[]{ll}k:popRank_{k,T}=I+1-debtRank_{i,T}&\mbox{ if }debtRank_{i,T}>I-K+1\\ k:popRank_{k,T}=K&\mbox{ otherwise }\end{array}\right$. (8)
Thus, the least-popular chair is reserved for the highest-ranked player, and the next-least-popular chair is reserved for the next-highest-ranked player, and so forth, until the most-popular chair is reached. The lowest-ranked IK+1I-K+1 players all lose because they are all assigned to the most-popular chair.
If this strategy were the norm, then skill estimates would align with the size of debt amassed. All chairs are assigned, so any player in a caste universe who chooses a chair other than their assigned chair will lose (assuming other players do not simultaneously defect from the caste strategy).
@@ 3.2 The turn-taking strategy
The turn-taking norm is displayed in the “Recommended Move 2” column of Fig. 1, and is an opposite of caste: Player $i$ selects chair $k$ by counting through the chairs, from most popular to least, until the count matches their non-reversed debt-rank or there are no chairs left.
$s_{i,T}=\left\{\begin{array}[]{ll}k:popRank_{k,T}=K+1-debtRank_{i,T}&\mbox{ if }debtRank_{i,T}<K-1\\ k:popRank_{k,T}=1&\mbox{ otherwise }\end{array}\right$. (9)
Thus, each of the lowest-ranked players wins, and only the IK+1I-K+1 highest-ranked players lose (all assigned to the least popular chair). Since each repetition in a turn-taking universe is won by whichever players are “owed” the most, the debts between players decrease or flip with each repetition, thus shuffling the rankings. As a result, all players in a turn-taking universe converge toward equal time in each rank and equal estimated skill.
As with the caste norm, all chairs are assigned, so any player in a turn-taking universe who chooses a chair other than their assigned chair will lose (assuming other players do not simultaneously defect from turn-taking). Caste and turn-taking strategies are not unique to MAD Chairs—for example, they can optimize average win-rate in two-player Chicken [12].
@@ 3.3 Accounting for non-learners
Each of these strategies can be adapted to account for play against potential non-learners. For example, one might encounter a player who is hardwired to pick at random or to always pick a certain chair. Such non-learners are liable to waste whichever chair the caste or turn-taking norm reserves for them, and such waste would gradually bring down average win rates across the entire population. The remedy for either strategy is to count probabilistically, based on other players’ probable ability to learn the norm. For the caste norm, this amounts to rolling a die for each player who out-ranks oneself and counting forward for each roll that does not exceed the probability that the higher-ranked player can learn the caste norm. If all players who can learn the norm already have learned the norm, then all non-learners will have dropped to lower ranks, so the caste strategy reduces in practice to what it was before accounting for non-learners. If one always faces the same set of players (i.e., if the multi-agent system is closed), then the caste strategy further reduces in practice to always picking the chair one picked in the previous repetition (i.e., raw territoriality).